Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.50 coming 3 years from today. The dividend should grow rapidly - at a rate of 21% per year - during Years 4 and 5; but after Year 5, growth should be a constant 8% per year.

If the required return on Computech is 17%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.
Answer #1

what is the value of the stock today

=1.50/(1+17%)^3+(1.50*(1+21%)^1)/(1+17%)^4+(1.50*(1+21%)^2)/(1+17%)^5+((1.50*(1+21%)^2*(1+8%))/(17%-8%))/(1+17%)^5

=14.93
Answer #2

Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.75 coming 3 years from today. The dividend should grow rapidly - at a rate of 32% per year - during Years 4 and 5; but after Year 5, growth should be a constant 7% per year. 

If the required return on Computech is 12%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.
Answer #3

SOLUTION :


Do = D1 = D2 = 0

D3 = 1.50 ($)

D4 = 1.50*1.21 = 1.815 ($)

D5 = 1.815*1.21 = 2.19615 ($)

D6 = 2.19615*1.08 = 2.371842 ($)

Thereafter constant growth @8% .


So, 

P5 = D6 /(r - g) = 2.371842/(0.17 - 0.08) = 26.3538 ($)


r = 17% = 0.17 

=> 1 + r = 1.17


Hence,


Current price 

= PV of future cash inflows 

= D3/1.17^3 + D4/1.17^4 + D5/1.17^5 + P5/1.17^5

= 1.815/1.17^3 + 2.19615/1.17^4 + 2.371842/1.17^5 + 26.3538/1.17^5

= 15.41 ($) (ANSWER).

answered by: Tulsiram Garg